The frequency of the allele A in a population at genetic equilibrium is 0.8. What is the expected frequency of the Aa genotype?
Answer #1

Frequency of Allele A is 0.8

p= 0.8, q= 1-0.8.= 0.2

Expected frequency of genotype Aa = 2pq = 2*0.8*0.2. =0.32